# Theory of Mind Enhances Collective Intelligence

### v1

@@ 4.2 Model interpretation
We make three observations regarding these two scenarios. First, we note that the interactions in the second scenario are a hypergraph in the sense that if any one of the nodes were removed in Figure 2, b. no remaining agent receives any payoff, all agents are necessary contributors to any single agent receiving a utility for their actions [38]. Second, it is straightforward to compute the collective intelligence of both scenarios using Equation 1 and the three stochastic variables $\{x_{1}^{t},x_{2}^{t},x_{3}^{t}\}$ when $s^{t}\in\{-1,1\}$ is uniformly distributed: In the first scenario it is 0 bits and for the second scenario it is 1 bit. This is a simple illustration of how a single agent, knowing the BPC model of two other agents, who are both naïve to all other agents’ BPC models, can improve the collective intelligence of the group by manipulating the network of interactions between agents. This is analogous to the results discussed in Woolley [121] in which group performance, using a measure of collective intelligence, is improved by having agents who have a theory of mind. Third, unlike evolutionary processes, the process in the second scenario has a goal-directed agent who is psychologically informed about the internal states of the other agents and so is able to make plans contingent on different configurations of the interaction network. This implies that $A_{1}$ understands the preferences implied by the co-factors of Equation 8, ($R+c$) and Equation 9, ($R+x_{1}$), influencing the behaviour of the other two agents. These preferences are the hidden (cognitive) variables discussed in Section 2.2, i.e. the Preferences aspect of the BPC framework. In terms of levels of ToM, $A_{1}$ only needs to ascribe hidden states to the other two agents but not the ability for $A_{2}$ or $A_{3}$ to ascribe hidden states to $A_{1}$, for example the NE can be reached by $A_{2}$ and $A_{3}$ by observing the other agent’s behaviour. So the ToM of $A_{1}$ is of second order and $A_{2}$ and $A_{3}$ only need a first order ToM. In this sense a ToM allows $A_{1}$ to reconfigure their social networks to further their goals at much shorter time-scales than the evolutionary time-scale of ecological networks.
@@ 4.3 Our perspective in context
We have reviewed some of the recent advances in how people, as complex agents, with a vast space of specialised, context-dependent behaviours, are able to coordinate their activities. In particular, finding ways to recombine our individual competencies to produce, in a short period of time, the appropriate collective competencies is a combinatorically complex task. The central theme of this article is that having a causal (generative) model of another person’s behaviour that reflects the current environmental and social context helps us manage this complexity. This is one way in which our solid brains produce the necessary liquid social structures that quickly build collective solutions with novel collective competencies. In this section we bring these ideas together and summarise our view.
@@
The separation of time-scales also plays an important distinction between learning processes that use the same theoretical foundation. The mathematical description of the processes that underpin the evolution of species has been shown to be equivalent to the statistical learning process that underpins Bayesian learning in individual agents [17, 113, 114], for both biological and artificial agents. What Bayesian learning provides at the level of the individual agent is an advantage in the speed of adaptation that Bayesian learning via evolution cannot achieve. Again, this suggests a universality of description that only varies qualitatively in terms of time-scales and the specifics of the mechanisms.

### v2

@@ 4.2 Model interpretation
We make three observations regarding these two scenarios. First, we note that the interactions in the second scenario are a hypergraph in the sense that if any one of the nodes were removed in Figure 2, b. no remaining agent receives any payoff, all agents are necessary contributors to any single agent receiving a utility for their actions [115]. Second, it is straightforward to compute the collective intelligence of both scenarios using Equation 1 and the three stochastic variables $\{x_{1}^{t},x_{2}^{t},x_{3}^{t}\}$ when $s^{t}\in\{-1,1\}$ is uniformly distributed: In the first scenario it is 0 bits and for the second scenario it is 1 bit. This is a simple illustration of how a single agent, knowing the BPC model of two other agents, who are both naïve to all other agents’ BPC models, can improve the collective intelligence of the group by manipulating the network of interactions between agents. This is analogous to the results discussed in Woolley [54] in which group performance, using a measure of collective intelligence, is improved by having agents who have a theory of mind. Third, unlike evolutionary processes, the process in the second scenario has a goal-directed agent who is psychologically informed about the internal states of the other agents and so is able to make plans contingent on different configurations of the interaction network. This implies that $A_{1}$ understands the preferences implied by the co-factors of Equation 8, ($R+c$) and Equation 9, ($R+x_{1}$), influencing the behaviour of the other two agents. These preferences are the hidden (cognitive) variables discussed in Section 2.2, i.e. the Preferences aspect of the BPC framework. In terms of levels of ToM, $A_{1}$ only needs to ascribe hidden states to the other two agents but not the ability for $A_{2}$ or $A_{3}$ to ascribe hidden states to $A_{1}$, for example the NE can be reached by $A_{2}$ and $A_{3}$ by observing the other agent’s behaviour. So the ToM of $A_{1}$ is of second order and $A_{2}$ and $A_{3}$ only need a first order ToM. In this sense a ToM allows $A_{1}$ to reconfigure their social networks to further their goals at much shorter time scales than the evolutionary time scale of ecological networks.
5 Strategic Inference, Theory of Mind, and the Dual Role of piKL
In strategic multi-agent environments, coordination and collaboration often depend not only on optimal decision-making, but on the agent’s ability to reason about the intentions, beliefs, and goals of others. This capability—often referred to as Theory of Mind (ToM)—becomes even more powerful when paired with language as a medium for influencing and inferring mental states. In this section, we develop a formal model that integrates reinforcement learning, ToM, and communication into a unified agent architecture. We also clarify two distinct interpretations of the piKL algorithm introduced in the Cicero AI framework.
5.1 Baseline Reinforcement Learning Policy
Let $s\in\mathcal{S}$ denote the environment state and $a\in\mathcal{A}$ the action selected by agent $i$. The baseline policy derived from reinforcement learning (RL) maximizes expected return:
$\pi^{\text{RL}}(a\mid s)=\arg\max_{a}Q(s,a)$
where $Q(s,a)$ is the value function learned via standard temporal-difference or policy-gradient methods.
5.2 Theory of Mind and Noisy Communication Channels
Each agent $i$ is assumed to have an internal, latent cognitive state $\theta_{i}\in\Theta$, encoding beliefs, goals, or preferences. Suppose agent $i$ communicates with agent $j$ via a message $m\in\mathcal{M}$ drawn from a noisy channel:
$P(m\mid\theta_{i})$
The receiving agent $j$ updates its belief over $\theta_{i}$ using Bayes’ rule:
$b_{j}(\theta_{i}\mid m)=\frac{P(m\mid\theta_{i})\cdot P(\theta_{i})}{\sum_{\theta^{\prime}_{i}}P(m\mid\theta^{\prime}_{i})\cdot P(\theta^{\prime}_{i})}$
This belief over the sender’s internal state allows agent $j$ to modulate its action policy accordingly.
5.3 ToM-Enhanced Policy via Belief Inference
The ToM-enhanced policy for agent $j$ is defined by taking the expectation over the inferred mental states of agent $i$:
$\pi_{j}^{\text{ToM}}(a\mid s,m)=\mathbb{E}_{\theta_{i}\sim b_{j}(\theta_{i}\mid m)}\left[\pi_{j}(a\mid s,\theta_{i})\right]$
This formulation permits the agent to adaptively align, cooperate, or compete based on its belief about the intentions or preferences of the other agent.
5.4 Strategic Pragmatics and Message Selection
Recognizing that the receiver is performing this inference, the speaker agent $i$ may choose its message strategically to manipulate beliefs:
$m^{*}=\arg\max_{m}\mathbb{E}_{\theta_{j}\sim b_{i}(\theta_{j})}\left[U_{i}(\theta_{j}(m))\right]$
This recursive, belief-aware messaging strategy is analogous to pragmatic language models such as the Rational Speech Acts (RSA) framework.
5.5 Two Interpretations of piKL
The piKL method measures divergence between policies. We describe two distinct formulations:
1. Anchor-Regularised piKL (Cicero)
In the Cicero system, the agent’s current policy $\pi_{i}$ is constrained by a divergence from an anchor policy $\pi_{\text{anchor}}$, learned from human gameplay:
$D_{\text{piKL}}^{\text{anchor}}=\mathrm{KL}\left[\pi_{i}(a\mid s)\parallel\pi_{\text{anchor}}(a\mid s)\right]$
This encourages the agent to remain near human-like, interpretable behavior. The total objective becomes:
$\mathcal{L}_{\text{anchor}}=\mathbb{E}_{(s,a)\sim\pi_{i}}\left[R(s,a)\right]-\lambda_{1}\cdot D_{\text{piKL}}^{\text{anchor}}$
2. ToM-Induced piKL
Alternatively, we define a divergence between the baseline RL policy and the ToM-enhanced policy:
$D_{\text{piKL}}^{\text{ToM}}=\mathrm{KL}\left[\pi^{\text{RL}}(a\mid s)\parallel\pi^{\text{ToM}}(a\mid s,m)\right]$
This measures how much the agent’s behavior is changed by incorporating belief inference and communicative reasoning.
5.6 Unified Objective
We combine both interpretations in a unified loss:
$\mathcal{L}_{\text{total}}=\mathbb{E}_{(s,a)\sim\pi_{i}}\left[R(s,a)\right]-\lambda_{1}\cdot\mathrm{KL}\left[\pi_{i}\parallel\pi_{\text{anchor}}\right]-\lambda_{2}\cdot\mathrm{KL}\left[\pi^{\text{RL}}\parallel\pi^{\text{ToM}}\right]$
Here, $\lambda_{1}$ and $\lambda_{2}$ control the influence of human-alignment and social-inference regularisation, respectively. Together, they allow an agent to balance exploitation, interpretability, and strategic communicative inference.
5.7 Interpretation
This model formalizes a cognitive architecture in which:
  • Reinforcement learning drives raw goal-seeking behavior;
  • Theory of Mind allows adaptation to others’ latent beliefs;
  • Language serves as a noisy yet informative conduit of cognitive intent;
  • Pragmatic reasoning aligns communicative goals with strategic planning;
  • Dual regularisation (via piKL) ensures the agent balances optimality with trustworthiness and social intelligence.
@@ 5.8 Our perspective in context
We have reviewed some of the recent advances in how people, as complex agents, with a vast space of specialised, context-dependent behaviours, are able to coordinate their activities. In particular, finding ways to recombine our individual competencies to produce, in a short period of time, the appropriate collective competencies is a combinatorically complex task. The central theme of this article is that having a causal (generative [118, 119]) model of another person’s behaviour that reflects the current environmental and social context helps us manage this complexity. This is one way in which our solid brains produce the necessary liquid social structures that quickly build collective solutions with novel collective competencies. In this section we bring these ideas together and summarise our view.
@@
The separation of time scales also plays an important distinction between learning processes that use the same theoretical foundation. The mathematical description of the processes that underpin the evolution of species has been shown to be equivalent to the statistical learning process that underpins Bayesian learning in individual agents [123, 124, 125], for both biological and artificial agents. What Bayesian learning provides at the level of the individual agent is an advantage in the speed of adaptation that Bayesian learning via evolution cannot achieve. Again, this suggests a universality of description that only varies qualitatively in terms of time scales and the specifics of the mechanisms.
The models we presented highlight the roles ToM plays in human social collectives such as deciding which relationships to develop and how these relationships aid in the collective processing of information towards an end goal. Being able to understand, predict, and manipulate these information-carrying connections is valuable in improving group dynamics (in human groups, AI groups, or hybrid human-AI groups) and for the future design of ToM-based AI. For example, knowing that ToM plays a role in group performance [54], quantifying and modelling (ToM-mediated) CI will shed light on the types of interventions or adjustments that can modulate it. Additionally, including AI agents with sophisticated ToM in human groups has the potential to ‘construct social niches’ (see below) more effectively, such as by modifying the incentives involved in a social setting (making some actions more desirable than others; cf. mechanism design [126]), or the connections between agents [127], or by simply mediating interactions with an increased capacity for memory, attention, and reasoning [128]. Recent efforts have leveraged ToM and AI for improving team effectiveness [129, 128] and understanding the link between ToM and deception [130], or ToM and expectation formation in markets [131]. Others suggest AI with ToM could assist humans in improving their ToM and communications skills [132], in negotiation, education (adaptation to students needs and knowledge), and games [133]; healthcare (to tailor care to patients’ mental states), self-driving cars (anticipating the actions of other cars), workplaces (employee mood and stress), and marketing [134].
To successfully apply ToM-based AI to these areas, we need to consider how we could define the appropriate incentives for an AI agent to adapt to a given set of social circumstances in a desirable manner. Our examples, while minimal, offer a conceptual framing for thinking about this human-AI integration. For example, the use of information theory to quantify the information processing provides a measure that can be used in an optimisation objective for an AI agent within a collective, incentivising an AI to act in ways that are constructive for the collective as a whole.
The difficulties of implementing socially aware algorithms in an AI are also made more complicated by human variations in expectations. For example, the famous trolley problem has often been used as an example of a moral dilemma that an autonomous AI that is in control of a car may confront during an emergency. To understand the complexities of this issue for an AI, Awad et al [135] collected  40 million individual decisions from 233 countries and territories in 10 languages. The questions focused on trolley-like moral dilemmas to examine cross-cultural ethical variation, showing three major clusters of countries and that these variations reflect differences in modern institutions and deep cultural traits. These results illustrate that our morality is not universal, there is a rich variety of expectations for the same moral dilemmas, and that any AI would need to reflect the local cultural variations in which they are used. In the Discussion next we cover the broader issue of how an AI will need to learn and adapt in order to fit the social circumstances the AI will be expected to be constrained by, very much like other socio-ecological systems.
To date the study of the collective intelligence in hybrid AI-human systems is in its infancy, however we already have many of the necessary tools to begin these investigations. ToM is experimentally well established in psychology and work has begun developing and implementing algorithms for an AI-ToM, see for example Mao et al’s recent review of empirical work on beliefs, desires and preferences for AI with ToM [136] as well as the computational, algorithmic, and experimental articles discussed in the Introduction above. What is not as well understood are the dynamics that agential AI will introduce into our human social ecology, and to that extent the study of machine behaviour [36] and new measures of collective intelligence as we have described here need to be applied to AI-human hybrid systems to better understand the collective outcomes, rather than focusing on the algorithmic foundations.